# Class-Conditional Distribution Balancing for Group Robust Classification

### v2

@@ IV Experimental Results
IV-A Experiments on Synthetic Data
In this section, we use synthetic datasets (i.e., cMNIST, cCIFAR10) to explore the effectiveness of our method in handling extreme bias under challenging multi-class settings. For each dataset, we control the extent of spurious correlations by varying the ratio of bias-conflicting samples in the training data, and test the performance on completely unbiased test data. Following [8], the bias-conflicting ratios are set to $\{0.5\%,1\%,2\%,5\%\}$ for both datasets, with $0.5\%$ representing an extremely biased scenario. The results are reported in Table III. From which, we have the following observations: ($i$) On cMNIST, our method consistently achieves the best performance, surpassing the second best with a large margin, even though some competing methods utilize bias annotations during model training or selection; ($ii$) On cCIFAR10, our method performs the best on average and significantly outperforms the others when the bias-conflicting ratio is as low as $0.5\%$. Considering that uLA employs resource-consuming MOCOV2+ [33, 69] pretraining and broad hyperparameter searching, our method provides a favorable trade-off between performance and computational complexity.
IV-B Experiments on Real-world Data
Experiments on real-world datasets, Waterbirds, CelebA, and CivilComments are conducted to test the effectiveness of our method in dealing with spurious correlations in practical scenarios. The results are reported in Table IV. With the supervision of bias annotations, GroupDRO demonstrates strong generalization performance on the worst group. On all three datasets, our method achieves competitive or superior performance compared to those utilizing bias information during validation. Despite employing a simpler learning process, our method performs comparably to uLA on CelebA, and significantly surpasses uLA and XRM on Waterbirds and CivilComments. By referring to Table III, we observe that bias supervision is more effective for simple binary classification tasks, e.g., GroupDRO achieves SOTA performance on Waterbirds and CelebA. However, for more complex multi-class classification tasks, bias-supervised methods fall significantly behind bias-free approaches(e.g., uLA and our CCDB).
To demonstrate that standard average accuracy on the validation set serves as an effective model selection criterion for CCDB, we record its training process on four datasets (cCIFAR10, Waterbirds, CelebA, and CivilComments), as shown in Figure 2. The average and worst-group accuracy curves on the validation set exhibit similar trends throughout the training process, with their optimal values occurring in regions where the test worst-group performance remains relatively high. This alignment represents an additional advantage of our method.
IV-C Ablation Study
Figure 4: Ablation study on the two key factors of CCDB. Left: the proportion $\gamma$ split from the training data; right: the threshold $T$ for sample weight clipping.
The proportion $\gamma$ split from the training data for biased feature extractor learning. By varying $\gamma$ in $\{10\%,30\%,50\%,70\%,100\%\}$, we visualize the t-SNE of the learned feature space during stage1 in Figure 3 and report the worst group accuracy of stage 3 in Figure 4 (left). In Figure 3, for cMNIST, as $\gamma$ decreases, the feature space exhibits an increasingly pronounced clustering effect aligned with bias annotations, indicating a feature space dominated by spurious factors. For Waterbirds, the dominance of spurious factors initially increases but then diminishes when $\gamma=0.1$, likely due to the smaller scale of the training data. Accordingly, with an appropriately small $\gamma\in[0.3,0.5]$, spurious factors are more effectively identified and balanced, leading to a classifier that relies more on causal factors and exhibiting improved generalization capabilities, see the optimal test accuracy (marked with a red star) in Figure 4 (left) for verification.
The threshold $T$ for sample weight clipping. The threshold $T$ controls the extent to which minority groups are emphasized, with larger values assigning greater weight to samples from minority groups. We evaluate the impact of $T$ by varying it from $0.5$ to $3.0$ and report the classification results in Figure 4(right). For datasets that are relatively less biased, e.g., Waterbirds, a threshold around $1.5$ yields the best performance. In contrast, for highly biased datasets, e.g., cCIFAR10 with a bias-conflicting ratio below $5\%$, a higher threshold around $2.5$ proves more effective.
@@ IV-C Ablation Study
Figure 5: The effect of intra-class compactness on the latent space. t-SNE visualization of the latent space learned on: (a) Waterbirds colored by group annotations; and (b) cCIFAR10 Colored by the spurious noise type attribute. With intra-class compactness, the distribution demonstrate a strong alignment with spurious information.
@@
The effect of intra-class compactness. The intra-class compactness constraint is applied to amplify bias attributes and smooth the feature space, thereby enhancing the model’s generalization ability. In this section, we conduct an ablation study on cMNIST, cCIFAR10, Waterbirds, and CelebA to evaluate the impact of intra-class compactness on the final performance of CCDB.
The results are presented in Figures 5 and 6. In Figure 5, under naive ERM training (without the intra-class compactness constraint), the feature distribution is skewed and primarily aligned with class information. In contrast, when intra-class compactness is applied, the distribution shifts to align more with bias attributes (such as background or noise type). Additionally, the structure of each cluster more closely resembles a Gaussian distribution compared to the results without intra-class compactness. Figure 6 shows the average and worst group accuracy with and without the intra-class compactness constraint. We observe that, for large-scale (e.g., CelebA) and simpler datasets (e.g., cMNIST), the constraint has a relatively modest effect on the final performance. However, for small-scale (e.g., Waterbirds) and complex (e.g., cCIFAR10) datasets, the constraint significantly improves performance, with gains of up to $16\%$ in accuracy.
IV-D Analysis on Sample Weights Learned by CCDB
Figure 7: The distribution of the learned sample weights within each group on four datasets. Samples from majority groups are given low weight, while those from minority groups are given high weight.
Figure 8: Activation visualizations of ERM and CCDB attention for Waterbirds and CelebA samples. Compared to ERM, CCDB more effectively focuses on the causal factors (foreground birds and hair color).
Figure 9: Boxplot of the feature correlations with bias and class information, before and after CCDB sample reweighting. Left: cMNIST; right: cCIFAR10.
To show that the proposed sample reweighting technique effectively reduces the correlation between spurious factors $\boldsymbol{z}_{\text{s}}$ and class labels $y$, we analyze the distribution of the learned sample weights and their impact on the spuriousness of the features extracted for classification.
Figure 7 shows the histograms of sample weights obtained from CelebA, Waterbirds, cCIFAR10 and cMNIST via our CCDB. The distributions are normalized within each group. For CelebA and Waterbirds, the weights of minority groups (man-blond, waterbird-land, landbird-water) flatten to values approximately ten times higher than those of the majority groups, which is consistent with our primal goal of up-weighting the groups with fewer training samples.
For cCIFAR10 and cMNIST, we divide the data into two subsets for demonstration: the bias-aligned subset where $\boldsymbol{z}_{\text{s}}$ is aligned with $y$, and the bias-conflicting subset where $\boldsymbol{z}_{\text{s}}$ conflicts with $y$. Obviously, the weights for the minority conflicting subset are concentrated around relatively high values, while the aligned subset has weights close to zero.
Figure 9 shows the effect of sample reweighting on the spuriousness of the learned features. We measure spuriousness by calculating the Pearson correlation coefficient [70] between each feature dimension and class/bias information, and present a boxplot of these coefficients. Note, the coefficients are computed on the test data, where no spurious correlation exists, meaning that class labels are solely correlated with causal factors and independent of spurious ones. Before applying sample reweighting, most extracted feature dimensions from ERM are dominated by spurious factors, exhibiting high correlation with bias information and low correlation with class labels (blue box). In contrast, after CCDB reweighting, the extracted features are predominantly causal, with strong correlation to class labels and weak correlation to bias information (yellow box). These observations highlight the effectiveness of CCDB in mitigating spurious correlations, thus creating a de-biased data distribution for downstream classification tasks.
To further demonstrate the effectiveness of the proposed CCDB in getting rid of spurious correlations, we compare the classifier’s attention before and after sample reweighting in Figure 8. The attention is visualized using xGradCAM [71]. Without sample reweighting, the model obtained from ERM primarily focuses on spurious factors, such as background or facial details. In contrast, after applying CCDB reweighting, the model is enforced to explore causal factors, e.g., the foreground birds and hair color, leading to more robust predictions.
TABLE V: Computation resource required in Stage2 on four datasets with a single GPU A40.
@@ IV-D Analysis on Sample Weights Learned by CCDB
IV-E The computational cost of stage 2
For optimization of Equation 7, we use Adam with learning rate $0.01$. Table V shows the storage and running time on all datasets. Obviously, the optimization is overall fast within a few minutes.
V Conclusion
Acknowledgments
This work was partly supported by the Talent Introduction Program of Dalian University of Technology, China (No. DUT23RC(3)041).

### v3

@@ 7 Viewing CCDB from a causal inference perspective
8 Data
We conduct experiments on five widely used benchmark datasets: cMNIST [46], cCIFAR10 [47], Waterbirds [21], CelebA [15], and CivilComments [48] to demonstrate the effectiveness of our method. The detailed statistics of each dataset are summarized in Tab. 3. Among these datasets, cMNIST and cCIFAR10 are ten-way classification tasks in which each class is dominated by a specific color or noise pattern, with varying bias ratios. Waterbirds, CelebA, and CivilComments are real-world datasets: Waterbirds and CelebA are image classification datasets where each class is strongly correlated with a particular background or gender attribute, respectively; CivilComments is a text classification dataset where the label is spuriously correlated with mentions of certain demographic identities.
9 Training setup
To make a fair comparison, we apply a 3-hidden layer MLP for cMNIST, a ResNet18 for cCIFAR10, a ResNet50 for Waterbirds and CelebA, and a BERT for CivilComments following existing methods [7, 18, 22]. Both ResNet18 and ResNet 50 are pretrained on ImageNet-1K, BERT is pretrained on Book Corpus and English Wikipedia. The proposed CCDB method consists of three stages, each optimized independently. The first two training stages span 20 epochs and 1000 iterations across all datasets (except CivilComments with 10 epochs); the third stage involves 2000 iterations for Waterbirds and CelebA, 5000 iterations for cMNIST and cCIFAR10, and 10 epochs for CivilComments. For Stage1, we use the Adam optimizer with a learning rate of $2\times 10^{-4}$ and default momentum parameters $\beta=(0.9,0.999)$ for all image datasets and AdamW optimizer with a learning rate of $1\times 10^{-5}$ for the NLP dataset; For stage2, the Adam optimizer with a learning rate of $1\times 10^{-2}$ is adopted across all datasets; The learning rate, along with weight decay and batch size used in stage3 for each dataset are detailed in Tab. 4.
On data preprocessing, we apply simple augmentations to each dataset as listed in Tab. 4, following [49]. These augmentations are specifically designed not to induce invariance to variations in the bias attribute. Consequently, the improvements reported in the main paper can be attributed to the proposed CCDB methodology, rather than to the use of bias-nullifying data augmentations.
All the results are averaged over 5 random seeds, with the best-performing model (highest validation accuracy) selected for each seed. From the results in Fig. 7, we conclude that CCDB can use either standard average accuracy or worst-class accuracy on a class-balanced validation set for model selection.
@@ 9 Training setup
Figure 8: The effect of intra-class compactness on the latent space. t-SNE visualization of the latent space learned on: (a) Waterbirds colored by group annotations; and (b) cCIFAR10 Colored by the spurious noise type attribute. With intra-class compactness, the distribution demonstrates a strong alignment with spurious information.
Figure 9: Boxplot of the feature correlations with bias and class information, before and after CCDB sample reweighting.
@@
10 Additional experimental results
10.1 On model selection for CCDB
To demonstrate that standard average accuracy on the validation set serves as an effective model selection criterion for CCDB, we record its training process on four datasets (cCIFAR10, Waterbirds, CelebA, and CivilComments), as shown in Fig. 7. The average and worst-group accuracy curves on the validation set exhibit similar trends throughout the training process, with their optimal values occurring in regions where the test worst-group performance remains relatively high. This alignment represents an additional advantage of our method.
10.2 The computational cost of stage 2
For optimization of Eq. 6, we use Adam with learning rate $0.01$. Tab. 5 shows the storage and running time on all datasets. Obviously, the optimization is overall fast within a few minutes.
Table 5: Computation resource required in Stage2 on four datasets with a single GPU A40.
@@ 10.2 The computational cost of stage 2
10.3 The effect of intra-class compactness
The intra-class compactness constraint is applied to amplify bias attributes and smooth the feature space, thereby enhancing the model’s generalization ability. In this section, we conduct an ablation study on cMNIST, cCIFAR10, Waterbirds, and CelebA to evaluate the impact of intra-class compactness on the final performance of CCDB.
The results are presented in Figs. 8 and 10. In Fig. 8, under naive ERM training (without the intra-class compactness constraint), the feature distribution is skewed and primarily aligned with class information. In contrast, when intra-class compactness is applied, the distribution shifts to align more with bias attributes (such as background or noise type). Additionally, the structure of each cluster more closely resembles a Gaussian distribution compared to the results without intra-class compactness. Fig. 10 shows the average and worst group accuracy with and without the intra-class compactness constraint. We observe that, for large-scale (e.g., CelebA) and simpler datasets (e.g., cMNIST), the constraint has a relatively modest effect on the final performance. However, for small-scale (e.g., Waterbirds) and complex (e.g., cCIFAR10) datasets, the constraint significantly improves performance, with gains of up to $16\%$ in accuracy.
10.4 Analysis on sample weights learned by CCDB
Fig. 9 shows the effect of sample reweighting on the spuriousness of the learned features. We measure spuriousness by calculating the Pearson correlation coefficient [51] between each feature dimension and class/bias information, and present a boxplot of these coefficients. Note, the coefficients are computed on the test data, where no spurious correlation exists, meaning that class labels are solely correlated with core factors and independent of spurious ones. Before applying sample reweighting, most extracted feature dimensions from ERM are dominated by spurious factors, exhibiting high correlation with bias information and low correlation with class labels (blue box). In contrast, after CCDB reweighting, the extracted features are predominantly core, with strong correlation to class labels and weak correlation to bias information (yellow box). These observations highlight the effectiveness of CCDB in mitigating spurious correlations, thus creating a de-biased data distribution for downstream classification tasks.